OpenStudy (anonymous):

So I have two more questions, the first one Alex has twice as much money as Jennifer. Jennifer has $6 less than Shannon. Together the have $54. How much does each have? The second one is There are three exams in a marking period. A student received 75 and an 81. What grade must the student earn on the last exam to get an average of no less than 80 for the marking period?

OpenStudy (anonymous):

Ok so lets lay it all out, to make things more clear... Let the money Alex has be X and let that of Jennifer be Y. We know that Alex has twice as much money as Jennifer, so X = 2Y. Let Shannon's money be Z. Since Jennifer has $6 less than Shannon, then: Z = Y + 6. And both, Jennifer and Shannon have $54 then Y + Y + 6 = 54 that is 2Y + 6 = 54. You can complete it on your own now...Ask for help if you get stuck with anything.

OpenStudy (anonymous):

So y =24 right?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

And X = 2Y = 2 * 24 = ? Z = Y + 6 = 24 + 6 = ? Remember, X is Alex's money, Y is Jennifer's and Z is Shannon's. So Alex has $48, Jennifer has $24 and Shannon has $30.

OpenStudy (anonymous):

Ah okay thanks. :) How about the second problem?

OpenStudy (pfenn1):

Well, the average of the three exams must be no less than 80 (or greater than or equal to 80 depending on how you look at it.) How would you go about finding the average of the scores on the three tests? Let x = the score of the third test?

OpenStudy (anonymous):

Ok...Lets get to the second problem. He got 75% on the first exam and 81% on the second one... As @pfenn1 stated, let X be the score of the third test...So the average will be (75 + 81 + X) / 3... And you need an average greater than or equal to 80% then the following equation does the work: \[\frac{ 75 + 81 + X }{ 3 } \ge 80\]

OpenStudy (anonymous):

So I just add 75 and 81 to get 156 then divide by 3 which will get me 52?

OpenStudy (pfenn1):

Nope, you need to rearrange the inequality to get x on one side and all the numbers on the other. So multiple both sides by 3 and subtract 75 + 81 or 156 from both sides to get x.

OpenStudy (anonymous):

So the answer is 84?

OpenStudy (anonymous):

The answer is \[X \ge 84\]

OpenStudy (pfenn1):

Actually, since you are solving an inequality, \[x \ge 3(80)-156\ge 84\]

OpenStudy (anonymous):

Ah thanks

OpenStudy (anonymous):

Try it...Like (84 + 75 + 81) / 3 and you'll get 80..anything lower than 84 will not yield the desired result...

OpenStudy (anonymous):

Well I got 163.33333...

OpenStudy (pfenn1):

?

OpenStudy (anonymous):

Oh wait nevermind I did it wrong

OpenStudy (anonymous):

Yep I got 80